Curriculum Vitae

Naoki Takebayashi

Institute of Arctic Biology &

Department of Biology and Wildlife 311 Irving I Bldg University of Alaska Fairbanks Fairbanks, AK 99775, USA Phone: (907) 474-1178 Fax: (907) 474-6967 e-mail: ntakebayashi@alaska.edu

Professional Experience

Associate Professor, University of Alaska, Fairbanks, September 2009 – current

Assistant Professor, University of Alaska, Fairbanks, October 2003 – current

Post-Doctoral Research Associate, Duke University, January 2001 - March 2003 Molecular evolution of self-incompatibility locus. (P.I. Dr. Marcy K. Uyenoyama) Development of statistical methods and computer programs to characterize evolutionary history of self-incompatibility locus from genealogy of S-alleles

Degrees

Ph.D. in Biology, Bloomington, Indiana, USA, December 2000 Dissertation title: Self-fertilization, genetic associations, and genetic architecture: understanding the components of mating-system evolution. Major: Evolutionary Biology, Minor: Mathematics Dissertation Adviser: Dr. Lynda F. Delph

B.S. in Biology, Kyushu University, Fukuoka, Japan, March 1991

Additional Education

Summer Institute in Statistical Genetics: Markov Chain Monte Carlo for genetics. North Carolina State University, Raleigh, NC, June 2002

Studied mathematical population genetics under Dr. Marcy K. Uyenoyama, Duke University. January - May, 1994.

Research Experience

Part-time Researcher - Kyushu Kankyou Kenkyusho, Japanese Governmental Facility (1990-1991)

Environmental Census; identified bird species and assessed their abundance.

Research Associate - Ono lab, Ecology Dept., Kyushu University, Fukuoka, Japan (Spring 1991)

Worked with Dr. Akihiro Yamane, helped trap feral cats to study their social structure.

Fellowships and Grants

NSF, Collaborative Research: Diversification of the self-incompatibility locus in the Papaveraceae, J. R. Kohn, N. Takebayashi, D. E. Wolf, 3/15/2007-3/14/2011, $227,019

NIH Alaska BRIN/INBRE New initiative (seed) internal grand, “Molecular evolution of self-incompatibility locus”, N. Takebayashi and D. E. Wolf, 7/1//2004–6/30/2005, $25,000

NIH Alaska BRIN/INBRE internal Equipment grand, “Molecular evolution of self-incompatibility locus: purchase of plant growth chamber”, N. Takebayashi and D. E. Wolf, 7/1//2004–12/31/2004, $20,000 Indiana Academy of Science Research Grant, November 1996 ($1100), May 1995 ($900), October 1995, ($500)

California Native Plant Society, Hardman Native Plant Research Award, November 1995, $300 (with Peter Morrell)

Sigma Xi, The Scientific Research Society, Grant-in-Aid of Research, March 1993, $400

Indiana University, Departmental Research Support Competition, 1993, 1995, 1996, 1997, 1998, 1999, $250 each year

Bayard Franklin Floyd Fellowship for Research in Plant Sciences, Indiana University, Summer 1993, 1994, 1995, 1996, 1997, 1998

Presentations

mtml-msBayes: Approximate Bayesian comparative phylogeography inference from multiple taxa and multiple loci with rate heterogeneity. Wen Huang, Naoki Takebayashi, Yan Qi, Michael Hickerson. Annual Meeting of the Society for Molecular Biology and Evolution. June 2010, Lyon, France

Molecular Evolution of S-locus in Alaskan Papaver species. Naoki Takebayashi, Takashi Miyake, Diana E. Wolf, Paul Heflinger, Vanessa Dominguez. Annual Meeting, Society for the Study of Evolution, June 2008, Minneapolis, MN

Breakdown of self-incompatibility in Arabidopsis lyrata ssp. kamchatica. Diana E. Wolf, Janette Steets, Dane Salter, Naoki Takebayashi. Annual Meeting of the Society for Molecular Biology and Evolution, June 2007, Halifax, Noca Scotia, Canada.

Molecular evolution of the S locus in Alaskan Arabidopsis lyrata ssp. kamchatica. Naoki Takebayashi, Diana Wolf, Dane Salter, Lacie Jo Wesbrook. Annual Meeting, Society for the Study of Evolution, June 2005, Fairbanks, AK.

Molecular evolution of an imprinted gene in Arabidopsis sp. Diana Wolf, Naoki Takebayashi, Dane Salter, Lacie Jo Wesbrook. Annual Meeting, Society for the Study of Evolution, June 2005, Fairbanks, AK.

Evolution of a polygenic selfing-rate modifier under inbreeding depression caused by a symmetric overdominant locus. Annual Meeting, Society for the Study of Evolution, June 2001, Knoxville, TN.

Association between floral traits and inbreeding depression in Gilia achilleifolia. Annual Meeting, Society for the Study of Evolution, June 1999, Madison, WI.

Association between floral traits and inbreeding depression in Gilia achilleifolia (Polemoniaceae), Poster. Annual Meeting, Society for the Study of Evolution, June 1997, Boulder, CO.

Is selfing an evolutionary dead-end? Phylogenetic relationships among populations with selfing rate variation in the California annual, Gilia achilleifolia (Polemoniaceae). Annual meeting, Society for the Study of Evolution, June 1996, St. Louis, MO.

Publications — Peer-reviewed

Wolf, D. E., J. .A. Steets, N. Takebayashi 2009. Selection for mixed mating by sexually transmitted diseases and other enemies. Evolution (submitted, in revision)

Jørgensen, M.H., A. K. Brysting, N. Takebayashi, J. A. Steet, J. J. Beecher, and D. E. Wolf, Recurrent origins of the amphi-Pacific allopolyploid Arabidopsis kamchatica (Brassicaceae). (submitted to Annals of Botany, in revision).

Paape, T., T. Miyake, N. Takebayashi, D. E. Wolf, J. R. Kohn, Evolutionary genetics of self-incompatibility in Papaveraceae. Genetics (in review).

Huang, W., N. Takebayashi, Y. Qi, M. J. Hickerson. 2011 MTML-msBayes: Approximate Bayesian compartive phylogeographic inference from multiple taxa and multiple loci with rate heterogeneity. BMC Bioinformatics 12:1-14 Steets, J. A., N. Takebayashi, J. M. Byrnes, D. E. Wolf, 2010. Heterogeneous selection on trichome production in

Alaskan Arabidopsis kamchatica. American Journal of Botany 97:1098-1108

Olson, M., A. Robertson, N. Takebayashi, S. Silim, W. Schroeder, P. Tiffin, 2010, Nucleotide Diversity and linkage disequilibrium in Balsam Poplar (Populus balsamifera), New Phytologist 186:526-536

Wood, T. E., N. Takebayashi, I. Mayrose, M. S. Barker, P. B. Greenspoon, L. H. Rieseberg 2009. The frequency of polyploid speciation in vascular plants. Proceedings of the National Academy of Sciences of the USA 106: 13875–13879

Miyake, T., N. Takebayashi, D. E. Wolf 2009. Possible diversifying selection in the imprinted gene, MEDEA, in Arabidopsis. Molecular Biology and Evolution 26: 843–857

Hickerson, M. J., E. Stahl, and N. Takebayashi 2007. msBayes: Pipeline for testing comparative phylogeographic histories using hierarchical approximate Bayesian computation. BMC Bioinformatics 8: Article Number 268

Takebayashi, N., D. E. Wolf and L. F. Delph 2006. Effect of variation in herkogamy on outcrossing within a population of Gilia achilleifolia. Heredity 96: 159-165.

Takebayashi, N., E. Newbigin, M. K. Uyenoyama 2004. Maximum-likelihood estimation of rates of recombination within mating type regions. Genetics 167: 2097–2109.

Uyenoyama, M. K. and N. Takebayashi 2004. A simple method for computing exact probabilities of mutation numbers. Theoretical Population Biology 65: 271–284.

Uyenoyama, M. K. and N. Takebayashi 2004. Genus-specific diversification of mating types. In The Evolution of Population Biology: Modern Synthesis, R. Singh and M. K. Uyenoyama (eds.). p. 254–271. Cambridge University Press, New York.

Wolf, D. E. and N. Takebayashi 2004. Pollen limitation and the evolution of androdioecy from dioecy. The American Naturalist 163: 122–137.

Hickerson, M. J., M. A. Gilchrist, and N. Takebayashi 2003. Calibrating a molecular clock from phylogeographic data, Moments and likelihood estimators. Evolution 57 (10): 2216-2225.

Takebayashi, N., P. B. Brewer, E. Newbigin, and M. K. Uyenoyama 2003. Patterns of variation within self-incompatibility loci. Molecular Biology and Evolution 20 (11): 1778-1794.

Wolf, D. E., N. Takebayashi and L. H. Rieseberg 2001. Predicting the risk of extinction through hybridization. Conservation Biology 15: 1039-1053

Takebayashi, N. and P. L. Morrell 2001. Is self-fertilization an evolutionary dead end? Revisiting an old hypothesis using a macroevolutionary approach. Invited Special Paper. American Journal of Botany 88: 1143–1150

Takebayashi, N. and L. F. Delph 2000. An association between a floral trait and inbreeding depression. Evolution 54: 840–846.

Other Publications

Miyake, T., N. Takebayashi, D. Salter, D. E. Wolf 2007. Natural selection on imprinted genes in outcrossing Arabidopsis lyrata and self-fertilizing A. thaliana. Genes and Genetic Systems 82: 543-543 (meeting abstract)

Miyake, T., N. Takebayashi, and D. E. Wolf 2007. Theoretical progress in evolution of self-incompatibility loci (title translated from Japanese). In Shokubutsu-no Shinka, K. K. Shimizu and M. Hasebe (eds.) p. 92–96. Shukan-sha, Tokyo, Japan

Teaching Experience

Mentoring:

two post-doctoral students, three graduate students (three current), eight undergraduates (one current) Statistical Phylogeography Workshop, American Museum of Natural History Southwestern Field Station (April,

2009, 2010, with Bryan Carstens and Mike Hickerson). Undergraduate level Introduction to Plant Biology (University of Alaska, Fairbanks) Graduate level Evolutionary Ecology of Plant Reproductive Biology (University of Alaska, Fairbanks) Graduate and undergraduate level Computational Biology (University of Alaska, Fairbanks) Graduate level Theoretical Population Genetics (University of Alaska, Fairbanks) Assistant instructor, Honors Molecular Biology (1997, Indiana University) Lab instructor, Microbiology laboratory for biology major (1998-1999, Indiana University) Lab instructor, Microbiology laboratory for Optometry students (1998-2000, Indiana University) Teacher certification in Japan for teaching high-school science (1991)

includes classwork and internship teaching at public junior high-school (Marugame West Junior High-school, Kagawa, Japan, 1990)

Technical Skills

Laboratory techniques: isozyme analysis with cellulose acetate and starch gels, DNA extraction, rtPCR, PCR,

cloning, DNA sequencing, flow cytometry Field and ecological methods Quantitative genetics: Experience with large-scale greenhouse breeding experiments including measurement of her-

itability and genetic correlations, and analysis with restricted maximum likelihood estimators Statistical analysis including SAS and S/R programming Computer modeling and programming in C, Fortran77, Perl, Bio-perl, Mathematica, and Unix Shells Facility with use and administration of Linux, OpenStep and Macintosh. Network administration based on TCP/IP (NFS, DHCP, LDAP, DNS, sendmail)

Service

Associate Editor of Molecular Biology and Evolution (Oct 2009-current) Reviewer for The American Naturalist, American Journal of Botany, Ecology, Evolution, Evolutionary Ecology, Ge-

netics, Molecular Ecology, Molecular Biology and Evolution, New Phytologist, and National Science Fundation (NSF).

Contributor to open-source software development such as R (a language and environment for statistical computing). Alaska Statewide High School Science Symposium judge and Woodriver elementary school Science Fair judge NIH Alaska INBRE, Management Advisory Committe member, faculty coordinator of Bioinformatics and Compu-

tational Biology

Memberships

Society for Molecular Biology and Evolution (2007–present) Society for the Study of Evolution (1996–present) The American Society of Plant Taxonomists (1996)

June, 2011